# Sample Selection Bias in Machine Learning for Healthcare

### v2

@@ 1 Introduction
SSB is well known for decades [27, 29, 3], and it has been discussed across several disciplines, including social science and econometrics [30, 31], environmental studies [67], finance [38], causality [34, 4], fairness [21], healthcare [48], and machine learning [71]. It is also an important consideration for clinical study design as the results of a biased study may not apply to a real patient population [6, 52]. Despite being well-known and a threat to the validity of algorithms [15], SSB has been scarcely studied in machine learning for healthcare [70]. This could be potentially because (i) often less proportion of patients are excluded from the study population, making it hard to find enough data to correct the bias, (ii) most of the algorithms are not developed for the real-world deployment, (iii) the bias is sometimes hidden under the assumed randomness in selection [55], (iv) risk estimation primarily belongs to causality, a principled approach to decision-making that examines cause-and-effect relationships. Causality has developed formal techniques, including causal diagrams, to address SSB [68, 57], and (v) SSB might not always lead to inaccurate predictions for the non-selected patients, as they might not influence the algorithm’s decision-making process.
@@
The rest of the paper is organized as: Section 2 provides a problem definition of SSB, while Section 3 presents a brief overview of literature related to SSB. The proposed research direction and T-Net and MT-Net are discussed in Section 4, followed by performance analysis with synthetic and semi-synthetic datasets in Section 5. Discussion on the significance of results is presented in Section 6, and finally, the concluding remarks are presented in Section 7.
@@ 2 Background
Suppose $\mathbf{x}\in\mathbf{X}$ is a $d$-dimensional feature vector representing a patient, $y\in\mathbf{y}$ is the corresponding outcome and $s\in\mathbf{s}$ is a random selection variable which takes value $1$ if patient $\mathbf{x}$ is selected in the study population or 00 otherwise. For the sake of clarity and simplicity, we focus on binary classification, i.e., risk prediction task. However, our work can be extended to other problem settings, such as regression. Let $\mathbf{X}=\mathbf{X}^{0}\cup\mathbf{X}^{1}\in\mathbb{R}^{d}$ be a random sample from the general population, where $\mathbf{x\in X^{1}}$ has corresponding label $y\in\mathbf{y^{1}}$ and are selected in the study population, i.e., $s=1$, and $\mathbf{x\in X^{0}}$ does not have access to corresponding label and are not selected into the study population, i.e., $s=0$. We assume access to labelled and unlabelled samples as $\mathbf{D^{1}}=\{(\mathbf{x}^{1}_{1},s_{1}=1,y^{1}_{1}),(\mathbf{x}^{1}_{2},s_{2}=1,y^{1}_{2}),...,(\mathbf{x}^{1}_{n},s_{n}=1,y^{1}_{n})\}$ of $n$ data points, and $\mathbf{D^{0}}=\{(\mathbf{x}^{0}_{1},s_{1}=0),(\mathbf{x}^{0}_{2},s_{2}=0),...,(\mathbf{x}^{0}_{m},s_{m}=0)\}$ of $m$ data points, respectively. If patients are selected non-randomly in the study population then $\mathbf{D^{1}}$ forms a biased distribution of true distribution $\mathbf{X}\times\mathbf{s}\times\mathbf{y}$, and leads to SSB. Thus, SSB has a distribution mismatch between the study population and the target population. Zadrozny, [71] first formalised SSB in machine learning and discussed four cases based on the dependence of $s$ on $(\mathbf{x},y)$, and also aligned these cases to missingness, treating SSB as a missingness handling issue, since the outcome is missing for the non-selected patients.
@@
In our study, we focus on the case (ii) since it is the most important SSB case in practice and it is more realistic to assume the case (ii) than to assume the case (i) [71]. Secondly, this is the most widely assumed case, while using multiple imputations to address SSB, in statistical and clinical literature [44]. SSB has connections to other machine learning problems, including missingness and covariate shift or domain adaption. However, SSB is a bit different to standard missingness handling problems where some attributes of a patient are missing while other patients from the same population have those attributes [12]. SSB is a special case of missingness problem where the outcome is missing rather than attributes [20], and it may be difficult to impute the missing labels of the non-selected patients as labels of the entire subpopulation are missing. SSB also intertwines with the broader concept of domain adaptation. Domain adaptation ensures machine learning algorithms trained on a source domain can perform well in a different target domain with varying feature distributions [42]. Specifically, covariate shift addresses differences in input feature distributions between source and target domains, i.e., a setting where $P^{\text{source}}(x)\neq P^{\text{target}}(x)$ while everything else remains unchanged. Since for SSB $P(s|\mathbf{x},y)=P(s|\mathbf{x})$ and $P(\mathbf{x}|s=1)\neq P(\mathbf{x}|s)$, this implies $P^{\text{study}}(\mathbf{x})\neq P^{\text{target}}(\mathbf{x})$. Therefore, SSB and covariate shift share fundamental similarities in addressing mismatches in feature distributions between the study population (source domain) and the target population (target domain). However, they originate from different settings because for SSB the target population is supposed to be a super-set of the study population while for covariate shift the target domain is supposed to be different than the source domain.
@@ 4 Methods
In this section, we present the proposed target population identification (TPI) approach – a new research direction – to address SSB in machine learning for healthcare for developing algorithms that utilise the available biased data to make accurate predictions. The key idea of our proposed approach is that we constrain the underlying machine learning algorithms, at deployment time, to a subpopulation seen by the algorithm during the training process, i.e., to a target subpopulation which is representative of the study population, without requiring to balance distributions of the two populations. This differs from the existing SSB correction approaches which forces alignment of the distributions of the study and the target populations, resulting in a loss of predictive performance of algorithms. Our novel approach is motivated by the observation that in healthcare the study and the target populations differ in terms of the non-selected patients. This is unlike domain adaptation where the target dataset is different to the source dataset and the aim is to perform better on the target dataset using information from the source dataset [42]. So for healthcare, an approach based on the identification of the target subpopulation representative of the study population can help in addressing SSB.
Figure 2: An illustration of the proposed target population identification (TPI) approach which employs a two-step evaluation process where first it identifies the target subpopulation representative of the study population, followed by predictions for the selected subpopulation.
@@ 5.2 Analysis with the Synthetic Data
Here, we present a comparative study of various techniques for handling SSB with the synthetic dataset. Since the performance could vary across different event rates, selection rates or dataset sizes, first we present a summary of the performance across different settings using boxplots in Fig. 4, and subsequent subsections study the effect of the selection/event rates and dataset sizes. We consider the event rates and the non-selection rates from [5, 10, 20, 30, 40]as we don’t expect large non-selection rates or event rates in real-world datasets, and dataset sizes from [1000, 2000, 3000, 4000, 5000] [^6] For a dataset size of 1000 with a non-selection rate and an event rate of 10% each, it implies there are 100 non-selected patients and 900 selected patients. 90 Of the 900 selected patients have positive outcomes while the rest 810 has negative outcomes., where each setting is repeated and averaged over 10 seeds. We observe large variations in the performance of all the techniques due to diverse settings, where extreme cases like low event rates, low non-selection rates or small dataset sizes could lead to a large drop in the performance.
@@ 5.3 Analysis with the Semi-synthetic Data
On analysing the performance of Naive technique as compared with Oracle, we observe an average drop of two percent and three percent for COVID-19 and Diabetes, respectively. Moreover, while Oracle performs the best in terms of median performance for the COVID-19 and Diabetes datasets, we observe some baselines performing close to Oracle. For example, in the Diabetes dataset, KLIEP and DANN show performance like Oracle. The small drop in average performance across different settings and the good performance of some baselines can be attributed to combination of two factors. First, except for Oracle and the proposed techniques, the remaining techniques make predictions for all patients (i.e., included and excluded). Second, for certain combinations of non-selection rates and event rates, the non-selected patients do not influence the decision function of the prediction model (This happens when the non-selected patients lie on the correct side of the decision boundary and away from the decision boundary that they do not have any role in deciding the decision boundary.). Consequently, all techniques that make predictions for all patients consistently make correct predictions for the non-selected patients, which improves their overall performance for all patients. This is also reflected in Fig. 7, where the maximum performance of some baselines is even higher than that of Oracle. MT-Net shows the least variation in the performance among all the techniques across all the settings. The smaller variation in the performance of MT-Net occurs because it has access to included as well as excluded data points, resulting in shared learning between the tasks, which is helpful with limited data settings. Similar to the synthetic dataset, Imputation shows the worst performance both in terms of mean performance as well as variation in the performance. Interestingly, the performance of KMM and Imputation for COVID-19, and MT-Naive, DANN and Imputation for Diabetes falls below Naive method which indicates that the bias handling negatively impacts the performance. We study the effect of the non-selection rates, event rates and dataset sizes in the following subsections.
@@ 7 Conclusion
Acknowledgements
This work was supported in part by the National Institute for Health Research (NIHR) Oxford Biomedical Research Centre (BRC) and in part by InnoHK Project Programme 3.2: Human Intelligence and AI Integration (HIAI) for the Prediction and Intervention of CVDs: Warning System at Hong Kong Centre for Cerebro-cardiovascular Health Engineering (COCHE). DAC was supported by an NIHR Research Professorship, an RAEng Research Chair, the InnoHK Hong Kong Centre for Cerebro-cardiovascular Health Engineering (COCHE), the NIHR Oxford Biomedical Research Centre (BRC), and the Pandemic Sciences Institute at the University of Oxford. GN is funded by NIHR (Grant number 302607) for a doctoral research fellowship. The views expressed are those of the authors and not necessarily those of the NHS, the NIHR, the Department of Health, the InnoHK – ITC, or the University of Oxford.
@@ Acknowledgements
The complete code used in the study will be released after acceptance at https://github.com/jmdvinodjmd/SSB_MLHC.

### v3

@@ 1 Introduction
SSB is well known for decades [28, 30, 3], and it has been discussed across several disciplines, including social science and econometrics [31, 32], environmental studies [68], finance [39], causality [35, 4], fairness [22], healthcare [49], and machine learning [72]. It is also an important consideration for clinical study design as the results of a biased study may not apply to a real patient population [6, 53]. Despite being well-known and a threat to the validity of algorithms [16], SSB has been scarcely studied in machine learning for healthcare [71]. This could be potentially because (i) often less proportion of patients are excluded from the study population, making it hard to find enough data to correct the bias, (ii) most of the algorithms are not developed for the real-world deployment, (iii) the bias is sometimes hidden under the assumed randomness in selection [56], (iv) risk estimation primarily belongs to causality, a principled approach to decision-making that examines cause-and-effect relationships [11]. Causality has developed formal techniques, including causal diagrams, to address SSB [69, 58], and (v) SSB might not always lead to inaccurate predictions for the non-selected patients, as they might not influence the algorithm’s decision-making process.
@@
The rest of the paper is organised as: Section 2 provides a problem definition of SSB, while Section 3 presents a brief overview of literature related to SSB. The proposed research direction and T-Net and MT-Net are discussed in Section 4, followed by performance analysis with synthetic and semi-synthetic datasets in Section 5. Discussion on the significance of results is presented in Section 6, and finally, the concluding remarks are presented in Section 7.
@@ 2 Background
Suppose $\mathbf{x}\in\mathbf{X}$ is a $d$-dimensional feature vector representing a patient, $y\in\mathbf{y}$ is the corresponding outcome and $s\in\mathbf{s}$ is a random selection variable which takes value $1$ if patient $\mathbf{x}$ is selected in the study population or 0 otherwise. For the sake of clarity and simplicity, we focus on binary classification, i.e., risk prediction task. However, our work can be extended to other problem settings, such as regression. Let $\mathbf{X}=\mathbf{X}^{0}\cup\mathbf{X}^{1}\in\mathbb{R}^{d}$ be a random sample from the general population, where $\mathbf{x\in X^{1}}$ has corresponding label $y\in\mathbf{y^{1}}$ and are selected in the study population, i.e., $s=1$, and $\mathbf{x\in X^{0}}$ does not have access to corresponding label and are not selected into the study population, i.e., $s=0$. We assume access to labelled and unlabelled samples as $\mathbf{D^{1}}=\{(\mathbf{x}^{1}_{1},s_{1}=1,y^{1}_{1}),(\mathbf{x}^{1}_{2},s_{2}=1,y^{1}_{2}),...,(\mathbf{x}^{1}_{n},s_{n}=1,y^{1}_{n})\}$ of $n$ data points, and $\mathbf{D^{0}}=\{(\mathbf{x}^{0}_{1},s_{1}=0),(\mathbf{x}^{0}_{2},s_{2}=0),...,(\mathbf{x}^{0}_{m},s_{m}=0)\}$ of $m$ data points, respectively. If patients are selected non-randomly in the study population then $\mathbf{D^{1}}$ forms a biased distribution of true distribution $\mathbf{X}\times\mathbf{s}\times\mathbf{y}$, and leads to SSB. Thus, SSB has a distribution mismatch between the study population and the target population. Zadrozny, [72] first formalised SSB in machine learning and discussed four cases based on the dependence of $s$ on $(\mathbf{x},y)$, and also aligned these cases to missingness, treating SSB as a missingness handling issue, since the outcome is missing for the non-selected patients.
@@
In our study, we focus on the case (ii) since it is the most important SSB case in practice and it is more realistic to assume the case (ii) than to assume the case (i) [72]. Secondly, this is the most widely assumed case, while using multiple imputations to address SSB, in statistical and clinical literature [45]. SSB has connections to other machine learning problems, including missingness and covariate shift or domain adaption. However, SSB is a bit different to standard missingness handling problems where some attributes of a patient are missing while other patients from the same population have those attributes [13]. SSB is a special case of the missingness problem where the outcome is missing rather than attributes [21], and it may be difficult to impute the missing labels of the non-selected patients as labels of the entire subpopulation are missing. SSB also intertwines with the broader concept of domain adaptation. Domain adaptation ensures machine learning algorithms trained on a source domain can perform well in a different target domain with varying feature distributions [43]. Specifically, covariate shift addresses differences in input feature distributions between source and target domains, i.e., a setting where $P^{\text{source}}(x)\neq P^{\text{target}}(x)$ while everything else remains unchanged. Since for SSB $P(s|\mathbf{x},y)=P(s|\mathbf{x})$ and $P(\mathbf{x}|s=1)\neq P(\mathbf{x}|s)$, this implies $P^{\text{study}}(\mathbf{x})\neq P^{\text{target}}(\mathbf{x})$. Therefore, SSB and covariate shift share fundamental similarities in addressing mismatches in feature distributions between the study population (source domain) and the target population (target domain). However, they originate from different settings because for SSB the target population is supposed to be a super-set of the study population while for covariate shift the target domain is supposed to be different than the source domain.
@@ 4 Methods
In this section, we present the proposed target population identification (TPI) approach – a new research direction – to address SSB in machine learning for healthcare for developing algorithms that utilise the available biased data to make accurate predictions. The key idea of our proposed approach is that we constrain the underlying machine learning algorithms, at deployment time, to a subpopulation seen by the algorithm during the training process, i.e., to a target subpopulation which is representative of the study population, without requiring to balance distributions of the two populations. The non-selected patients are referred to a clinician, ensuring they receive appropriate care while maintaining the reliability of algorithmic predictions. This differs from the existing SSB correction approaches which forces alignment of the distributions of the study and the target populations, resulting in a loss of predictive performance of algorithms. Our novel approach is motivated by the observation that in healthcare the study and the target populations differ in terms of the non-selected patients. This is unlike domain adaptation where the target dataset is different to the source dataset and the aim is to perform better on the target dataset using information from the source dataset [43]. So for healthcare, an approach based on the identification of the target subpopulation representative of the study population can help in addressing SSB.
Figure 2: An illustration of the proposed target population identification (TPI) approach which employs a two-step evaluation process where first it identifies the target subpopulation representative of the study population, followed by predictions for the selected subpopulation. The non-selected patients are referred to a clinician.
@@ 5.2 Analysis with the Synthetic Data
Here, we present a comparative study of various techniques for handling SSB with the synthetic dataset. Since the performance could vary across different event rates, selection rates or dataset sizes, first we present a summary of the performance across different settings using boxplots in Fig. 4, and subsequent subsections study the effect of the selection/event rates and dataset sizes. We consider the event rates and the non-selection rates from [5, 10, 20, 30, 40]% as we don’t expect large non-selection rates or event rates in real-world datasets, and dataset sizes from [1000, 2000, 3000, 4000, 5000] [^6] For a dataset size of 1000 with a non-selection rate and an event rate of 10% each, it implies there are 100 non-selected patients and 900 selected patients. 90 Of the 900 selected patients have positive outcomes while the rest 810 has negative outcomes., where each setting is repeated and averaged over 10 seeds. We observe large variations in the performance of all the techniques due to diverse settings, where extreme cases like low event rates, low non-selection rates or small dataset sizes could lead to a large drop in the performance.
@@ 5.3 Analysis with the Semi-synthetic Data
On analysing the performance of Naive technique as compared with Oracle, we observe an average drop of two percent and three percent for COVID-19 and Diabetes, respectively. Moreover, while Oracle performs the best in terms of median performance for the COVID-19 and Diabetes datasets, we observe some baselines performing close to Oracle. For example, in the Diabetes dataset, KLIEP and DANN show performance like Oracle. The small drop in average performance across different settings and the good performance of some baselines can be attributed to combination of two factors. First, except for Oracle and the proposed techniques, the remaining techniques make predictions for all patients (i.e., included and excluded). Second, for certain combinations of non-selection rates and event rates, the non-selected patients do not influence the decision function of the prediction model (This happens when the non-selected patients lie on the correct side of the decision boundary and away from the decision boundary that they do not have any role in deciding the decision boundary.). Consequently, all techniques that make predictions for all patients consistently make correct predictions for the non-selected patients, which improves their overall performance for all patients. This is also reflected in Fig. 7, where the maximum performance of some baselines is even higher than that of Oracle.
MT-Net shows the least variation in the performance among all the techniques across all the settings. The smaller variation in the performance of MT-Net occurs because it has access to included as well as excluded data points, resulting in shared learning between the tasks, which is helpful with limited data settings. Similar to the synthetic dataset, Imputation shows the worst performance both in terms of mean performance as well as variation in the performance. Interestingly, the performance of KMM and Imputation for COVID-19, and MT-Naive, DANN and Imputation for Diabetes falls below Naive method which indicates that the bias handling negatively impacts the performance. We study the effect of the non-selection rates, event rates and dataset sizes in the following subsections.
@@ 7 Conclusion
The complete code is available at https://github.com/jmdvinodjmd/SSB_MLHC.
Acknowledgements
This work was supported in part by the National Institute for Health Research (NIHR) Oxford Biomedical Research Centre (BRC) and in part by the ITC InnoHK “Oxford-CityU Hong Kong Centre for Cerebrocardiovascular Health Engineering” (COCHE). DAC was supported by an NIHR Research Professorship, an RAEng Research Chair, and the Pandemic Sciences Institute at the University of Oxford. GN is funded by the NIHR (Grant number 302607) for a doctoral research fellowship. The views expressed are those of the authors and not necessarily those of the NHS, the NIHR, the Department of Health, the InnoHK – ITC, or the University of Oxford.
We also thank the reviewers for their insightful feedback, which has greatly improved the quality of this work.